OpenStudy (anonymous):

Find all solutions for (tan^2 θ − 16)(2 cos θ + 1) = 0

OpenStudy (campbell_st):

well the solutions are where \[\tan^2(\theta) - 16 = 0\] and \[2\cos(\theta) + 1 = 0\] hope it helps

OpenStudy (anonymous):

oh wow I kept thinking I had to substitute them within the equation and then factor. I'm an idiot, thanks.